can someone plase show me how to solve this problem
y=x+7
y=-x+2


y+3=-3(x-6)

Answer by Alan3354(30993)

can someone plase show me how to solve this problem
y=x+7
y=-x+2

y+3=-3(x-6) Is this supposed to be part of the problem?
------------------------
y= x+7
y=-x+2 Add the 2 eqns
2y = 9
y = 4.5
-------
Sub into eqn 1
4.5 = x + 7
x = -2.5
----------